MOLECULAR ORBITAL STUDY OF THE HYDRATED CLUSTERS OF STRONG ACIDS WITH WATER MOLECULES

\underline{S. RE}, Y. SUZUKI, AND Y. OSAMURA, Department of Chemistry, Faculty of Science, Rikkyo University , 3-34-1 Nishi-ikebukuro, Toshima-ku, Tokyo 171, Japan; H. F. SCHAEFER III, Center for Computational Quantum Chemistry, University of Georgia, Athens, GA 30602, USA.

The molecular structures and vibrational frequencies of the hydrated clusters of the strong acids, HCl•(H_{2}O)_{n} and H_{2}SO_{4}•(H_{2}O)_{n} ; n=1∼5, are examined by employing the density functional molecular orbital method. When the number of water molecules is small, the hydrated clusters take the hydrogen-bonded structures without proton transfer. In the cases of n≥4, the proton transferred forms become dominant and there are many isomeric forms for both the direct ionpair X^{−}•H_{3}O^{+}•(H_{2}O)_{n−1}, and indirect ion-pair X^{−}•(H_{2}O)_{n−1}•H_{3}O^{+} structures, where X=Cl, HSO_{4}. The calculated IR spectra of the stable clusters clearly indicate the large red-shifts of the stretching frequencies of the H-Cl and the O-H bond of H_{2}SO_{4}. The difference of the hydrated clusters of HCl, H_{2}SO_{4}, and HF is also discussed.